OpenStudy (anonymous):

Let f(x)=5x-7 and g(x)=(x+4)^2. Find the composite function f(g(x))

OpenStudy (satan ):

Do you have to simplify it, or can you leave it un-simplified?

OpenStudy (anonymous):

I think it has to be simplified

OpenStudy (satan ):

Okay, start by simplifying g(x).

OpenStudy (satan ):

\(\Large (x+4)^2=(x+4)(x+4)\) FOIL it.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

is that all I have to do?

OpenStudy (satan ):

But after that, you have to plug whatever you get for \(g(x)\) into \(f(x)\), and distribute that.

OpenStudy (satan ):

So \(g(x)=x^2+8x+16\) would get plugged in: \(\Large f(g(x))=5(x^2+8x+16)-7\) Now you have to simplify that, and that will be your simplified composite function.

OpenStudy (satan ):

Oh, alternative form is \(\Large f(g(x))=5((x+4)^2)-7\)

OpenStudy (satan ):

Both are right, just depends on which one you want.

OpenStudy (anonymous):

ohh ok so I just simplify that and I get the answer?

OpenStudy (satan ):

That is correct.

OpenStudy (anonymous):

ok thank you!

OpenStudy (satan ):

Anytime :)